OpenStudy (anonymous):

equation for line tangent to the curve. x=-cost, y=8+sint, t=pi/2 and find the d2y/dx2

sam (.sam.):

Where's the Equation of curve?

OpenStudy (lgbasallote):

andthe equation of the curve is....

sam (.sam.):

dx/dt=sin(t) dy/dt=cos(t) =================== dy/dx=cost/sint dy/dx=cot (t) d2y/dx^2=-csc^(2)t d2y/dx^2=-csc^(2)(pi/2) d2y/dx^2=-1

sam (.sam.):

dy/dx=cot (t) dy/dx=cot (pi/2) ----------------- t=pi/2 x=-cost y=8+sint x=0 y=8 ----------------------- y-y1=m(x-x1) y-8=cot (pi/2) (x-0) y= x cot (pi/2) +8